OpenStudy (anonymous):

medal to the first one who will explain clearly ... how do you solve p ..... 3p+3q=p the answer is p=-3/2p , but need help understanding the steps ...

OpenStudy (anonymous):

You have it almost right...

OpenStudy (anonymous):

\[q=-2p/3\]

OpenStudy (anonymous):

i got that at first too but got it wrong the answer was P=-3/2p so i dont get it

OpenStudy (anonymous):

You have to divide both sides by a constant to simplify the equation, so you'd come up with p+q=p/3

OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

just a sec ill sent you the question

OpenStudy (anonymous):

solve for p. 3(p+q)=P A) q=-2/3p B) q=-3/2p C) p=-2/3q D) p=-3/2q

OpenStudy (anonymous):

the correct answer was D

OpenStudy (anonymous):

ok thx ill ask the teacher about this question .

OpenStudy (anonymous):

i got it